OpenStudy (anonymous):

wut is the derivative of 2te^-t

OpenStudy (anonymous):

2t(-e^-t)+2e^-t

OpenStudy (callisto):

d/dt 2te^-t = (e^-t) d/dt (2t) + (2t) d/dt (e^-t) = 2e^-t + (2t) (e^-t) (-1) = 2e^-t - 2t (e^-t) = 2e^-t (1-t)

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome :)